# Can We Treat Noisy Labels as Accurate?

### v1

@@ Abstract
Noisy labels significantly hinder the accuracy and generalization of machine learning models, particularly due to ambiguous instance features. Traditional techniques that attempt to correct noisy labels directly, such as those using transition matrices, often fail to address the inherent complexities of the problem sufficiently. In this paper, we introduce EchoAlign, a transformative paradigm shift in learning from noisy labels. Instead of focusing on label correction, EchoAlign treats noisy labels ($\tilde{Y}$) as accurate and modifies corresponding instance features ($X$) to achieve better alignment with $\tilde{Y}$. EchoAlign’s core components are (1) EchoMod: Employing controllable generative models, EchoMod precisely modifies instances while maintaining their intrinsic characteristics and ensuring alignment with the noisy labels. (2) EchoSelect: Instance modification inevitably introduces distribution shifts between training and test sets. EchoSelect maintains a significant portion of clean original instances to mitigate these shifts. It leverages the distinct feature similarity distributions between original and modified instances as a robust tool for accurate sample selection. This integrated approach yields remarkable results. In environments with 30% instance-dependent noise, even at 99% selection accuracy, EchoSelect retains nearly twice the number of samples compared to the previous best method. Notably, on three datasets, EchoAlign surpasses previous state-of-the-art techniques with a substantial improvement.
@@ 1 Introduction
The rapid advancement of neural networks highlights the importance of learning from noisy labels (LNL) (Tan and Le, 2019; Dosovitskiy et al., 2021; Stiennon et al., 2020; Chen et al., 2023a). Though cost-effective for large datasets, web crawling, and crowdsourcing can introduce noisy labels that hinder machine learning model generalization (Yu et al., 2018; Li et al., 2017; Welinder et al., 2010; Zhang et al., 2017; Natarajan et al., 2013; Gu et al., 2023). For example, recent studies indicate this label noise in pretraining data impacts out-of-distribution generalization of foundation models on downstream tasks (Chen et al., 2023a, 2024). Noisy labels are categorized as random, class-dependent, or instance-dependent, with the latter two being particularly challenging due to ambiguous instance features. This ambiguity complicates the distinction between mislabeled examples and those belonging to other classes (Menon et al., 2018; Xia et al., 2020; Yao et al., 2023a; Bai et al., 2023).
Prior work has largely approached LNL from either noise-modeling-free or noise-modeling standpoints. The noise-modeling-free methods, such as extracting examples with small losses (Han et al., 2018; Yu et al., 2019; Wang et al., 2019), can only do clean sample selection and cannot correct noisy labels to clean labels, which waste much supervision information. On the other hand, noise-modeling approaches explicitly consider the label-noise generation process (Scott et al., 2013; Scott, 2015; Goldberger and Ben-Reuven, 2016). These methods operate on the principle that the clean class posterior can be deduced using the noisy class posterior, derived from noisy data, in conjunction with a transition matrix (Berthon et al., 2021). The concept is that, given a transition matrix, an optimal classifier mirroring clean data’s behavior can be trained solely using sufficient noisy data (Reed et al., 2014; Liu and Tao, 2015). Nonetheless, accurately modeling the general process, specifically the transition matrix, is ill-posed by only exploiting noisy data (Xia et al., 2019; Cheng et al., 2020). The additional assumptions required for this model are often hard to validate and may not hold in real-world datasets, as indicated by recent studies (Xia et al., 2020; Yao et al., 2023b; Liu et al., 2023). This mismatch frequently results in an inability to eliminate label noise thoroughly.
@@
In this paper, we challenge the conventional approach to label noise with instance modification. Instead of attempting to correct noisy labels, we strategically adjust instances to better align with their labels, even if those labels are incorrect. This groundbreaking new direction of LNL is visually depicted in Figure 1. At its core, instance modification tackles the underlying cause of label noise. Drawing on principles from causal learning (Neuberg, 2003; Peters et al., 2017; Yao et al., 2021), we view the generation of instance-dependent label noise through a causal lens, as illustrated in the causal graph of Figure 2. Here, the observed noisy labels ($\tilde{Y}$) emerge from the instances ($X$), influenced by latent variables ($Z$). In crowdsourcing scenarios, the difficulty of labeling blurred or ambiguous instances often leads to errors. Instead of trying to guess the ‘true’ label in these cases, modifying the instance itself to make it more easily classifiable can be more effective. Therefore, instance modification offers a practical solution for particularly challenging instances that are difficult to label correctly – those that are inherently ambiguous for both humans and machines.
While instance modification offers a breakthrough, it presents significant challenges at both the instance and dataset levels (More discussions can be found in the Appendix E). At the instance level, a critical challenge lies in maintaining the instance’s core characteristics. Modifying instances, particularly those with incorrect labels, to strictly align with potentially unrelated labels can lead to substantial characteristic shifts. Imagine modifying a dog image to resemble an airplane – the resulting image might be unnatural or distorted, compromising the dog’s intrinsic characteristics like posture, contour, and depth information. Meanwhile, it’s also crucial to preserve these intrinsic characteristics during modification for instances with correct labels. At the dataset level, instance modification can induce covariate shifts between training and test sets (Han et al., 2022a, b). These covariate shifts can hinder a model’s ability to generalize effectively to unseen real-world data.
@@
To tackle these challenges, we propose a simple yet effective framework EchoAlign (§4). EchoAlign has two key components: EchoMod and EchoSelect. EchoMod modifies instances using controllable generative models to align them with noisy labels while preserving their intrinsic characteristics. By controlling the image generation process using the original image and noisy label as guiding information, EchoMod ensures the overall quality and consistency of the training data. EchoSelect addresses the issue of covariate shift by strategically selecting original instances with correct labels to balance the distribution between original and modified instances in the training data. This selection leverages a novel insight: after instance modification, the Cosine feature similarity distribution between original and modified images shows a clear distinction between clean and noisy samples. EchoSelect employs this insight through a unique similarity metric to select the most reliable training instances to reconstruct the learnable set for further supervised or self-supervised training.
Our contributions and the key findings are summarized as follows: (1) We introduce a transformative change in how noisy labels are approached. Instead of focusing on correcting them, we advocate treating noisy labels as accurate and modifying instances to align with them. This paradigm shift is supported by our theoretical analysis (§3). (2) We present EchoAlign, a framework specifically designed to address the challenges of instance modification. It employs EchoMod for controlled instance modification and EchoSelect for strategic sample selection (§4). (3) We empirically validate the advantages of instance modification and demonstrate EchoAlign’s exceptional performance in noisy environments. Across three datasets with diverse noise types, EchoAlign consistently outperforms state-of-the-art techniques, achieving significant accuracy gains (§5).
@@ 2 Related Work
Learning with Noisy Labels  Research in this domain has predominantly followed two paths: (1) Noise-modeling-free methods: These methods primarily rely on the memorization effects observed in deep neural networks (DNNs). They capitalize on the tendency of DNNs to prioritize learning simpler (clean) examples before memorizing more complex (noisy) ones (Arpit et al., 2017; Wu et al., 2020; Kim et al., 2021). Techniques include early stopping (Han et al., 2018; Nguyen et al., 2020; Liu et al., 2020; Xia et al., 2021; Lu et al., 2022; Bai et al., 2021), pseudo-labeling (Tanaka et al., 2018), and leveraging Gaussian Mixture Models in a semi-supervised learning context (Li et al., 2020). (2) Noise-modeling methods: This category of approaches focuses on estimating a noise transition matrix, which models how clean labels might be corrupted into noisy observations. However, accurately modeling the noise process is challenging when relying solely on noisy data (Xia et al., 2019; Cheng et al., 2020). Existing studies rely on assumptions that might not hold in real-world datasets (Xia et al., 2020; Yao et al., 2023b; Liu et al., 2023). Consequently, these methods often struggle to fully eliminate label noise, especially when dealing with structured dependent noise patterns, such as subclass-dominant label noise (Bai et al., 2023).
Generative Models  Recent advances in generative models, including variational auto-encoders, generative adversarial networks, and diffusion models, have transformed applications with their exceptional sample generation capabilities (Du et al., 2023; Wang et al., 2023; Franceschi et al., 2023). Diffusion models, particularly known for their superior output control, delicately denoise signals (Zhang et al., 2023; Kingma et al., 2021). While these models hold promise for noisy label scenarios, existing approaches like Dynamics-Enhanced Generative Models (DyGen) (Zhuang et al., 2023) and Label-Retrieval-Augmented Diffusion Models (Chen et al., 2023b) still focus on enhancing predictions or retrieving latent clean labels. Our work introduces a fundamentally different approach. We adopt controllable generative models, treating noisy labels as correct and aligning instances to them. This bypasses the challenges of traditional noise modeling and focuses on improving the training data itself. Controllable generative models are a powerful subset of generative models that allow for precise control over the generated outputs, such as ControlNet (Zhang et al., 2023), iPromptDiff (Chen et al., 2023c). Unlike traditional generative models, which generate images based on random noise, controllable generative models take control information (e.g., text descriptions, class labels, or reference images) as input (Bose et al., 2022). This control information guides the generation process, ensuring that the generated outputs align with the desired characteristics.
@@ 3 Analysis
Problem Definition  In addressing the challenges posed by learning from noisy labels (LNL), we formally define the problem and introduce the concept of instance modification within a mathematical framework. Let $\mathcal{X}$ represent the input space of instances and $\mathcal{Y}$ the space of labels. In the traditional LNL setting, each instance $X\in\mathcal{X}$ is associated with a noisy label $\tilde{Y}\in\mathcal{Y}$, which may differ from the true label $Y\in\mathcal{Y}$. The goal is to learn a mapping $f:\mathcal{X}\rightarrow\mathcal{Y}$ that predicts the true label $Y$ as accurately as possible, despite the presence of noisy labels. Instance modification diverges from the conventional approach of directly correcting noisy labels $\tilde{Y}$ to match the true labels $Y$. Instead, it proposes adjusting each instance $X$ to better align with its given noisy label $\tilde{Y}$.
@@
Mathematically, this involves transforming each instance $X$ into a modified instance $X^{\prime}$, such that $f(X^{\prime})$ more closely approximates $\tilde{Y}$, leveraging the inherent information contained within the noisy label itself.
Theoretical Analysis  According to the causal learning framework (Liu et al., 2023; Yao et al., 2021), the noise can often be represented as a function of both the instance features and external factors, encapsulated by latent variables $Z$. We assume the causal relations (commonly occurring in crowdsourcing scenarios) are represented in the causal graph as illustrated in Figure 2, where $Z$ represents latent variables that affect both $X$ and $\tilde{Y}$ indirectly through $X$. Instance modification aims to transform $X$ into $X^{\prime}$ such that the modified instance $X^{\prime}$ aligns better with $\tilde{Y}$ under the assumption that $\tilde{Y}$ contains some information about the true label $Y$. Accordingly, we can deduce the effectiveness of instance modification as follows.
@@
Assume that the noisy labels are generated by a stochastic process influenced by latent variables $Z$, where $\tilde{Y}=h(Y,Z)$ and $Y$ are the true labels. Let $T$ be a transformation such that $X^{\prime}=T(X,\tilde{Y};\theta)$, where $\theta$ is chosen to optimize the alignment of $X^{\prime}$ with $\tilde{Y}$. Then, under this transformation, the predictive performance of a model trained on $(X^{\prime},\tilde{Y})$ is theoretically improved compared to a model trained on $(X,\tilde{Y})$ in terms of:
@@
    Alignment: The mutual information between $X^{\prime}$ and $\tilde{Y}$, $I(X^{\prime};\tilde{Y})$, is maximized relative to $I(X;\tilde{Y})$, indicating better alignment of modified instances with their noisy labels.
@@
    Generalization: Modifications in $X^{\prime}$ lead to better generalization. By transforming the original instances to better align with their noisy labels, the model trained on $X^{\prime}$ is less to overfit to the noise and more capable of capturing the true underlying patterns in the data.
This improvement is contingent upon the assumption that the noise model $h$ and the transformation $T$ are appropriately defined and that the latent variable model adequately captures the underlying causal structure of the data. More details and proofs can be found in the Appendix A.
This theoretical framework 3.1 suggests that instance modification, by aligning more closely with noisy but informative labels $\tilde{Y}$, can leverage the inherent structure and causality in the data to enhance learning. It demonstrates that instance modification improves alignment between instances and noisy labels, reduces information loss, and ultimately leads to better generalization. These insights provide several key motivations for the design of our method. Firstly, the improvements in alignment highlight the importance of modifying instances to embed noisy label information directly. This motivates the use of controllable generative models in EchoAlign, which can effectively incorporate label information into the instance features. Secondly, ensuring a minimal distribution difference between $X$ and $X^{\prime}$ is crucial. EchoMod generates $X^{\prime}$ with small distribution differences from $X$, while EchoSelect retains clean samples to control distribution differences, ensuring better generalization on test data.
@@
Thirdly, the improvement in estimation stability indicates that using modified features can result in more consistent and reliable model predictions, driving the focus on preserving the essential characteristics of the data during transformation to reduce variability and enhance stability.
Analyzing Feature Similarity Distributions  In this study, we address the challenges of instance modification which can induce distribution shifts between the training and test sets. Preserving clean original instances is crucial to mitigating these shifts. Existing sample selection methods (e.g., small loss (Han et al., 2018)) often falter under complex label noise conditions, such as instance-dependent noise, necessitating a more precise selection strategy. To this end, we find an interesting phenomenon: Clean samples generally exhibit higher similarity between features of original and modified images, indicating minimal semantic and label changes after modification, whereas noisy samples display lower similarity due to significant semantic and label adjustments. Utilizing the feature similarity distributions between original and modified instances emerges as a robust tool for enhancing sample selection accuracy. These distinctions are visually represented in Figure 3. The similarity is computed using the CLIP ViT-B-32 feature extractor (Radford et al., 2021) on the CIFAR-10 dataset with 30% instance-dependent noise. We use ControlNet (Zhang et al., 2023) to modify instances. The black dashed line indicates the sample threshold achievable by the previous best method at 96% accuracy (Yang et al., 2022). In contrast, EchoSelect, at 96% accuracy, can retain the samples in the yellow section. In environments with 30% instance-dependent noise, EchoSelect retains nearly twice the number of samples at 99% accuracy. Statistical validation using the Kolmogorov-Smirnov test confirmed significant differences in the distributions (p-value < 0.001), demonstrating the utility of feature similarity as a robust metric for identifying clean samples within noisy datasets.

### v2

# Can We Treat Noisy Labels of Ambiguous Instances as Accurate?

@@ Abstract
Noisy labels significantly hinder the accuracy and generalization of machine learning models, particularly when resulting from ambiguous instance features that complicate correct labeling. Traditional approaches, such as those relying on transition matrices for label correction, often struggle to effectively resolve such ambiguity, due to their inability to capture complex relationships between instances and noisy labels. In this paper, we propose EchoAlign, a paradigm shift in learning from noisy labels. Unlike previous methods that attempt to correct labels, EchoAlign treats noisy labels ($\tilde{Y}$) as accurate and modifies corresponding instances ($X$) to better align with these labels. The EchoAlign framework comprises two main components: (1) EchoMod leverages controllable generative models to selectively modify instance features, achieving alignment with noisy labels while preserving intrinsic instance characteristics such as shape, texture, and semantic identity. (2) EchoSelect mitigates distribution shifts introduced by instance modifications by strategically retaining a substantial subset of original instances with correct labels. Specifically, EchoSelect exploits feature similarity distributions between original and modified instances to accurately distinguish between correctly and incorrectly labeled samples. Extensive experiments across three benchmark datasets demonstrate that EchoAlign significantly outperforms state-of-the-art methods, particularly in high-noise environments, achieving superior accuracy and robustness. Notably, under 30% instance-dependent noise, EchoSelect retains nearly twice the number of correctly labeled samples compared to previous methods, maintaining 99% selection accuracy, thereby clearly illustrating the effectiveness of EchoAlign. The implementation of EchoAlign is publicly available at https://github.com/KevinCarpricorn/EchoAlign/tree/main.
keywords:
Learning from Noisy Labels , Controllable Generative Models , Instance Modification , Feature Alignment , Sample Selection , Robust Machine Learning.
[…]
[label1]organization=Sydney AI Centre, The University of Sydney,city=Sydney, postcode=2050, state=NSW, country=Australia
[…]
[label2]organization=Electrical and Mathematical Sciences and Engineering Division, King Abdullah University of Science and Technology,city=Thuwal, postcode=23955, country=Saudi Arabia
[…]
[label3]organization=School of Software, Shandong University,city=Jinan, postcode=250100, country=China
@@ 1 Introduction
The rapid advancement of neural networks has underscored the significance of learning from noisy labels (LNL)  (Tan and Le,, 2019; Dosovitskiy et al.,, 2021; Stiennon et al.,, 2020; Chen et al., 2023a, ). Although web crawling and crowdsourcing provide cost-effective means for collecting large datasets, they often introduce noisy labels that hinder model generalization (Yu et al., 2018b, ; Li et al.,, 2017; Welinder et al.,, 2010; Zhang et al.,, 2017; Natarajan et al.,, 2013; Gu et al.,, 2023). Recent studies have highlighted that label noise in pretraining data adversely affects the out-of-distribution generalization of foundation models in downstream tasks (Chen et al., 2023a, ; Chen et al.,, 2024). Noisy labels are generally categorized as random, class-dependent, or instance-dependent, with the latter two posing particular challenges due to ambiguous instance features, making it difficult to distinguish mislabeled examples from true class instances (Menon et al.,, 2018; Xia et al.,, 2020; Yao et al., 2023a, ; Bai et al.,, 2023).
Prior research has primarily approached LNL through either noise-modeling-free or noise-modeling frameworks. Noise-modeling-free techniques, such as filtering out high-loss examples (Han et al.,, 2018; Yu et al.,, 2019; Wang et al.,, 2019), are limited to selecting clean samples and do not address the potential for correcting incorrect labels, thereby discarding valuable supervisory information. In contrast, noise-modeling approaches explicitly consider the label-noise generation process (Scott et al.,, 2013; Scott,, 2015; Goldberger and Ben-Reuven,, 2016), often employing a transition matrix to relate noisy labels to their clean counterparts (Berthon et al.,, 2021). Theoretically, an optimal classifier can be trained with sufficient noisy data and an accurate transition matrix (Reed et al.,, 2014; Liu and Tao,, 2015). However, estimating this matrix is inherently ill-posed due to uncertainty and variability in noisy data (Xia et al.,, 2019; Cheng et al.,, 2020). Moreover, these models often rely on additional assumptions, such as the exact nature of the noise, which are challenging to validate and may not hold in real-world datasets, leading to suboptimal performance (Xia et al.,, 2020; Yao et al., 2023b, ; Liu et al.,, 2023). Traditional label correction methods are particularly limited when dealing with label noise caused by ambiguous features. For example, in datasets collected through web crawling, an image labeled as ‘dog’ might actually depict a cartoon or a product featuring a dog, which makes label correction challenging and often impractical.
@@
In this paper, we introduce a novel perspective on handling label noise by employing instance modification rather than correcting labels. Instead of attempting to correct noisy labels, we adjust instances to better align with their labels, even if those labels are incorrect. This innovative approach, illustrated in Figure 1, directly addresses the root cause of label noise. Leveraging causal learning principles (Neuberg,, 2003; Peters et al.,, 2017; Yao et al.,, 2021), we model instance-dependent label noise from a causal perspective, as depicted in the causal graph of Figure 3. Specifically, we consider how different factors, such as instance characteristics and latent variables, contribute to the generation of noisy labels, enabling us to better understand and address the root causes of label noise. In crowdsourcing scenarios, for instance, ambiguous or blurred instances are more prone to labeling errors. Instead of attempting to infer the ‘true’ label, modifying the instance to make it more distinguishable can be more effective. For example, in medical imaging, if a tumor is labeled as malignant but its visual features are too subtle, enhancing the image to highlight relevant features can assist both models and humans in identifying it more accurately (Mirza et al.,, 2023). Similarly, in sentiment analysis, modifying ambiguous sentences to be more explicit can better align them with their intended sentiment labels, thereby reducing ambiguity and improving classification accuracy.
(a) Characteristic Shift.
However, instance modification presents challenges at both the instance and dataset levels. At the instance level, a key challenge is modifying instances while preserving their essential characteristics, such as shape, texture, or color patterns. These features are crucial for distinguishing related categories. Excessive alterations can distort these defining features, leading to a phenomenon we refer to as the characteristic shift (Figure 2(a)). For example, when a wolf is mislabeled as a dog, transforming the wolf image into a typical dog might eliminate important shared features such as body shape and texture, which are critical for distinguishing between wolves and dogs. At the dataset level, instance modification may introduce a distribution shift (Figures 2(b) and 2(c)), where the statistical distribution of modified instances deviates from that of the original instances, potentially affecting model generalization to real-world data (Han et al., 2022a, ; Han et al., 2022b, ). Empirical observations, as visualized through T-SNE projections, reveal that modified instances may occupy a distinct feature space from their unaltered counterparts, altering the training dynamics. Addressing these shifts requires a balanced framework that preserves essential characteristics while effectively managing distribution differences between original and modified instances to ensure robustness in real-world applications.
@@
To address these challenges, we propose a simple yet effective framework, EchoAlign (§4). EchoAlign consists of two key components: EchoMod and EchoSelect. EchoMod modifies instances using controllable generative models, ensuring alignment with noisy labels while preserving intrinsic characteristics. EchoSelect mitigates covariate shifts by selecting original instances with correct labels, maintaining a balanced distribution between original and modified data. This selection is guided by a novel insight: after instance modification, the cosine feature similarity between original and modified images reveals distinctions between correctly and incorrectly labeled samples. EchoSelect uses this similarity metric to curate a reliable training set, improving robustness and accuracy in both supervised and self-supervised training.
Our key contributions and findings are summarized as follows:
  1. 1.
    We introduce a transformative shift in addressing label noise by modifying instances to align with noisy labels instead of correcting them, supported by theoretical analysis (§3);
  2. 2.
    We present EchoAlign, a framework featuring EchoMod for controlled instance modification and EchoSelect for strategic sample selection (§4);
  3. 3.
    We empirically validate the benefits of instance modification and demonstrate EchoAlign’s superior performance in noisy environments across three types of noisy data and real-world scenarios, significantly outperforming state-of-the-art methods in accuracy (§5).
@@ 2 Related Work
Learning with Noisy Labels  Research in this domain has predominantly followed two paths: (1) Noise-modeling-free methods: These methods primarily rely on the memorization effects observed in deep neural networks (DNNs), which tend to learn simpler (clean) examples before memorizing more complex (noisy) ones (Arpit et al.,, 2017; Wu et al.,, 2020; Kim et al.,, 2021). Techniques include early stopping (Han et al.,, 2018; Nguyen et al.,, 2020; Liu et al.,, 2020; Xia et al.,, 2021; Lu et al.,, 2022; Bai et al.,, 2021), pseudo-labeling (Tanaka et al.,, 2018), and leveraging Gaussian Mixture Models in a semi-supervised learning context (Li et al.,, 2020). (2) Noise-modeling methods: These approaches focus on estimating a noise transition matrix, modeling how clean labels can become corrupted into noisy observations. However, accurately modeling this noise process is particularly challenging when relying solely on noisy data (Xia et al.,, 2019; Cheng et al.,, 2020). Many existing studies depend on assumptions that may not hold in real-world datasets (Xia et al.,, 2020; Yao et al., 2023b, ; Liu et al.,, 2023). Consequently, these methods often struggle to effectively handle structured noise patterns, such as subclass-dominant label noise (Bai et al.,, 2023).
Generative Models  Recent advances in generative models, including variational auto-encoders, generative adversarial networks, and diffusion models, have transformed applications with their exceptional sample generation capabilities (Du et al.,, 2023; Wang et al.,, 2023; Franceschi et al.,, 2023). Diffusion models, known for their superior output control, are particularly effective at denoising signals (Zhang et al.,, 2023; Kingma et al.,, 2021). While these models hold promise for noisy label scenarios, existing approaches like Dynamics-Enhanced Generative Models (DyGen) (Zhuang et al.,, 2023) and Label-Retrieval-Augmented Diffusion Models (Chen et al., 2023b, ) still focus primarily on enhancing predictions or retrieving latent clean labels. Our work takes a fundamentally different approach. We leverage controllable generative models, treating noisy labels as correct and aligning instances with these labels, thus bypassing the challenges of traditional noise modeling and focusing on improving the quality of training data. Controllable generative models, such as ControlNet (Zhang et al.,, 2023) and iPromptDiff (Chen et al., 2023c, ), enable precise control over the generated outputs. Unlike traditional generative models which generate images from random noise, controllable generative models use control information (e.g., text descriptions, class labels, or reference images) as input (Bose et al.,, 2022), guiding the generation process to ensure that the outputs align with the desired characteristics.
@@ 3 Analysis
Problem Definition In addressing the challenges posed by learning from noisy labels (LNL), we formally define the problem and introduce the concept of instance modification within a mathematical framework. Let $\mathcal{X}$ represent the input space of instances and $\mathcal{Y}$ the space of labels. In the traditional LNL setting, each instance $X\in\mathcal{X}$ is associated with a noisy label $\tilde{Y}\in\mathcal{Y}$, which may differ from the true label $Y\in\mathcal{Y}$. The goal is to learn a mapping $f:\mathcal{X}\rightarrow\mathcal{Y}$ that predicts the true label $Y$ as accurately as possible, despite the presence of noisy labels. Instance modification diverges from the conventional approach of directly correcting noisy labels $\tilde{Y}$ to match the true labels $Y$. Instead, we propose adjusting each instance $X$ to better align with its given noisy label $\tilde{Y}$. Mathematically, this involves transforming each instance $X$ into a modified instance $X^{\prime}$, such that $f(X^{\prime})$ aligns more closely with $\tilde{Y}$, leveraging the inherent information contained within the noisy label itself.
Theoretical Analysis  According to the causal learning framework (Liu et al.,, 2023; Yao et al.,, 2021), the noise can often be represented as a function of both the instance features and external factors, encapsulated by latent variables $Z$. We assume that the causal relations (commonly occurring in crowdsourcing scenarios) are represented by the causal graph as illustrated in Figure 3, where $Z$ represents latent variables that affect both $X$ and $\tilde{Y}$ indirectly through $X$. Instance modification aims to transform $X$ into $X^{\prime}$ such that the modified instance $X^{\prime}$ better aligns with $\tilde{Y}$ under the assumption that $\tilde{Y}$ contains partial information about the true label $Y$. Accordingly, we can deduce the effectiveness of instance modification as follows.
@@
[…]
@@
[…]
@@
    Generalization: Modifications in $X^{\prime}$ lead to better generalization. By transforming the original instances to better align with their noisy labels, the model trained on $X^{\prime}$ is less likely to overfit to the noise and more capable of capturing the true underlying patterns in the data.
This improvement is contingent upon the assumption that the noise model $h$ and the transformation $T$ are appropriately defined and that the latent variable model adequately captures the underlying causal structure of the data. More details and proofs can be found in A.
Theorem 3.1 suggests that instance modification, by aligning more closely with noisy but informative labels $\tilde{Y}$, can leverage the inherent structure and causality in the data to enhance learning. It demonstrates that instance modification improves alignment between instances and noisy labels, reduces information loss, and ultimately leads to better generalization. These insights provide several key motivations for the design of our method. First, the improvements in alignment highlight the importance of modifying instances to embed noisy label information directly. This motivates the use of controllable generative models in EchoAlign, which can effectively incorporate label information into the instance features. Second, ensuring a minimal distribution difference between $X$ and $X^{\prime}$ is crucial. EchoMod generates $X^{\prime}$ with small distribution differences from $X$, while EchoSelect retains clean samples to control distribution differences, ensuring better generalization on test data. Third, the improvement in estimation stability indicates that using modified features can result in more consistent and reliable model predictions, motivating a focus on preserving the essential characteristics of the data during transformation to reduce variability and enhance both statistical stability and robustness in model performance.
Analyzing Feature Similarity Distributions  In this study, we address the challenges of instance modification, which can induce distribution shifts between the training and test sets. Preserving clean original instances is crucial to mitigating these shifts. Existing sample selection methods (e.g., small loss (Han et al.,, 2018)) often falter under complex label noise conditions, such as instance-dependent noise, necessitating a more precise selection strategy. To this end, we find an interesting phenomenon: Clean samples generally exhibit higher similarity between features of original and modified images, indicating minimal semantic and label changes after modification, whereas noisy samples display lower similarity due to significant semantic and label adjustments. Utilizing the feature similarity distributions between original and modified instances emerges as a robust tool for enhancing sample selection accuracy. These distinctions are visually represented in Figure 4.
@@
The similarity is computed using the CLIP ViT-B-32 feature extractor (Radford et al.,, 2021) on the CIFAR-10 dataset with 30% instance-dependent noise. We use ControlNet (Zhang et al.,, 2023) to modify instances. The black dashed line indicates the sample threshold achievable by the previous best method at 96% accuracy (Yang et al.,, 2022). In contrast, EchoSelect, at 96% accuracy, can retain the samples in the yellow section. In environments with 30% instance-dependent noise, EchoSelect retains nearly twice as many samples at 99% accuracy. Statistical validation using the Kolmogorov-Smirnov test confirmed significant differences in the distributions (p-value < 0.001), demonstrating the utility of feature similarity as a robust metric for identifying clean samples within noisy datasets.
[…]
(a) 50% symmetric noise
(b) 45% pairflip noise
(c) 50% instance-dependent noise